# Towards Privacy-aware Mental Health AI Models: Advances, Challenges, and Opportunities

### v1

@@ Introduction
Privacy evaluation tests synthetic data robustness against membership and attribute inference attacks [113, 44, 43, 42]. Membership inference attacks identify if an individual is part of the real dataset, with outliers being especially vulnerable. This is critical for mental health datasets, where diverse, small participant groups make outlier protection a priority. Metrics like privacy gain [44] and outlier similarity [43] are used for outlier privacy evaluation. Ensuring zero duplication of real sessions in synthetic data, measured through reproduction rate [43], is essential. Further memorization, overfitting and identification metrics include memorization coefficients [114] and $\epsilon$-identifiability [115]. Attribute inference attacks exploit known attributes to deduce sensitive details like mental health conditions. Additional metrics include distance-based metrics [43, 42].
@@
Synthetic data quality is assessed through faithfulness (similarity to real-world data) and diversity (lexical, semantic, and topic variation) [97, 46]. Faithfulness is measured via vocabulary overlap and semantic consistency[46]. However, these need to be supplemented by expert evaluations. Experts rate transcripts on naturalness, empathy, helpfulness, and safety [26]. Other metrics include four-level rating systems [116]. For multimodal data, ratings assess naturalness of speech, gestures, and their coherence and contextual plausibility [49, 50]. To reduce reliance on human evaluations, LLMs can automatically rate synthetic data on attributes like professionalism, comprehensiveness, authenticity and safety [47], or use psychological measures like the Working Alliance Inventory [46]. The utility is also evaluated through model performance on downstream tasks [48, 27] using synthetic data or synthetic data augmentation with real data. Ensuring synthetic data utility in mental health applications involves preserving relevant features while maintaining diversity and faithfulness.
@@
Privacy evaluation tests the robustness of trained models against malicious attacks. For language models, this includes measuring exposure through canary insertion and membership inference attack accuracy [32]. Another aspect is assessing whether model embeddings used for mental health predictions inadvertently reveal private attributes like location, age, gender, or identity [33, 110, 111]. Most privacy-aware training methods utilize DP-SGD [31], where the privacy guarantee is theoretically quantified by $\epsilon$-value [35, 34, 36, 37] (which determines the distance within which errors are considered to be zero in SGD). If FL is involved in the training process, its robustness against FL specific attacks [105, 106, 107] should be determined along with leakage through local model weights and communicated gradients [108]. Evaluating privacy through these methods ensures the robustness of training techniques.
@@
Figure 2: Our proposed pipeline for data collection and model training to develop privacy-aware mental health AI models.
@@
The first step in building mental health AI systems is collecting video recordings of therapy sessions. However, due to the sensitive nature of this data, there is a high risk of privacy breaches, including identification and impersonation attacks. To mitigate these risks, explicit, informed consent from patients is mandatory. The consent form should specify that the data will be anonymized, stored securely, and used exclusively for research purposes. Additionally, an ethics committee must review and approve the data collection and storage procedures to ensure compliance with privacy regulations and ethical standards. The audio recording should be performed using two channel recorder, such that therapist and patient voice can be untangled easily. The video recording should be focused on the patient showing their full face and posture so that facial expressions, gaze and body language can be studied. The recorded sessions should be transcribed by involved researchers or through local ASR systems to assure privacy.
@@
Once collected, the data must be anonymized or replaced with synthetic data to protect patient privacy. This decision is based on the dataset size and diversity. Training transformer models requires a large amount of diverse data for generalization. Therefore, if a dataset contains a small number of participants or less diversity among participants, synthetic data should be generated to improve the data utility. The generated data can also be augmented with real datasets for training models. In case the dataset already contains a large number of diverse data points, only data anonymization suffices. Anonymization of real data involves processing text, audio, and video to remove or replace PII. For text transcripts, LLMs can be employed to detect and redact PII [21]. However, their performance on small datasets may be limited, necessitating training on augmented datasets for improved generalization [77, 78, 20]. For audio recordings, PII can be replaced by synthesizing matching audio segments [22], and multi-speaker anonymization techniques can be used to disguise voices while preserving conversational dynamics [45]. Video recordings should undergo face anonymization using advanced methods such as FIVA [93]. Alternatively, synthetic data can be generated to ensure it does not relate to real individuals. This can be achieved using multimodal LLMs [118] capable of role-playing as therapists and patients [46]. Alternatively, these models can be used to generate realistic therapy sessions through zero-shot or few-shot prompting techniques [27], ensuring the generated data bears no resemblance to actual individuals.
@@
To ensure the efficacy and safety of the data, it is necessary to evaluate the privacy-utility trade-off after anonymization or synthetic data generation. Privacy evaluation of anonymized data includes testing re-identification risks using adversarial models [38], other related datasets [51] and measuring the effectiveness of techniques like multi-speaker anonymization through metrics such as EER [23, 24, 82, 84, 85] and FAR [45] in speaker verification. Voice anonymization should also be evaluated for unlinkability [85, 40] and robustness against attacks [23, 24, 82]. For video recordings, the privacy risks can be assessed using face verification systems and face reconstruction attacks[93] to determine the degree of obfuscation. Synthetic data must be rigorously tested against membership inference attacks and attribute inference attacks to ensure it does not inadvertently reveal details of the real dataset [113, 44, 43, 42]. Outlier leakage is another critical concern, especially in mental health datasets, where outliers are more prevalent due to the diversity and small size of participant groups. Metrics such as privacy gain [44], outlier similarity [43], and reproduction rate [43] are effective for evaluating these risks. These empirical privacy evaluations should be performed on different languages, cultures and demographics to obtain a more holistic idea about the privacy guarantees. Moreover, multimodal privacy measures need to be developed to understand and test cross-modal vulnerabilities. The empirical privacy measures should also be accompanied by theoretical guarantees similar to $\epsilon$-value in DP. Utility evaluation should assess the usefulness of the data for mental health diagnosis tasks, focusing on information preservation [112, 38], intonation preservation [23], conversational diversity [23, 82], naturalness [23, 45], and emotional feature retention [24, 92]. LLMs can be leveraged to automatically evaluate the utility of synthetically generated data on dimensions such as comprehensiveness, professionalism, authenticity, and safety [47] or psychological measures like the working alliance inventory [46].
@@
Privacy-preserving methods, particularly differential privacy, are crucial during this stage. Mental health models either consist of pre-trained models fine-tuned on mental health data or use pre-trained models to extract embeddings to train lightweight modality fusion layers for specific mental health diagnosis tasks. In the first approach, pre-trained models should use differential privacy-based fine-tuning methods [35, 34] to ensure privacy. In the second approach, fusion layers should be trained with DP-SGD [31] to ensure no privacy leakages from the trained layers. For even greater privacy Local Differential Privacy for Federated Learning (LDP-FL) [108, 109] can be used. This can be especially useful when datasets from different institutes are involved and they are required to be stored in the collected institutes for privacy.
@@
While Miao et al. [45] provided a benchmark for MSA, they assume weak attack models where the attacker does not have knowledge about the used anonymization scheme. However, in real-life situations, the attacker might possess knowledge of the anonymization strategy, and thus, privacy evaluation of MSA should be performed with stronger adversaries and subsequently develop better anonymization strategies. Moreover, overlapping segments in multi-speaker conversations like therapy sessions present another vulnerability that can be utilized by attackers. While Miao et al. [45] tested the ability of attackers to infer speakers from overlapping segments, in real-life situations, attackers might possess the ability to separate the speakers in overlapping segments and identify them. This shows the need to create stronger MSA schemes.
@@
Current video anonymization methods show vulnerability to leaking the gender and age of people even after face obfuscation [25]. Mental health datasets might contain very few participants from a certain gender or age group, thus leaking such private information could lead to identification. Recording the body language of patients could help in mental health diagnosis; however, it can also reveal the gender of the patient if only face anonymization is performed [96]. Moreover, current methods are also prone to demographic unfairness [41]. Thus it is essential to develop fair and improved video anonymization techniques that can prevent leakage of private information like age and gender.
@@
While we discuss various data anonymization processes for text, audio and video modalities, most of them do not provide any theoretical guarantees like DP provides in privacy-aware model training. In text modality, word-level or sentence-level perturbations through DP provide theoretical guarantees[79]. However, they significantly reduce the utility of the text [79], necessitating better anonymization techniques with privacy guarantees for text. For voice anonymization, Shamsabadi et al. [85] provided privacy guarantees in single-speaker anonymization settings. However, therapy sessions require multi-speaker anonymization, leaving room to create a multi-speaker anonymization algorithm with privacy guarantees. DP-based face anonymization techniques already exist for images [94]. For videos DP-based methods focus on making two objects within the video indistinguishable [95]. However, its direct applicability for preventing facial recognition in therapy videos is unclear. Thus the method needs to go through privacy and utility evaluations and we need to develop more specialized DP-based video face anonymization methods. In synthetic data generation DP-based methods for theoretical guarantees have been explored for text [99, 100], tabular data [119, 120], multimodal tabular and 3D image data [121] generation and with FL [122]. However, no DP-based methods have been developed for multimodal therapy session generation.
@@
Although significant progress has been made in anonymizing individual data modalities, such as text, audio, and video in isolation, there remains a lack of research on anonymization techniques for multimodal data. Multimodal datasets inherently carry cross-modal features, where information from one modality may inadvertently expose sensitive details from another. For instance, lip-reading from video can reveal PII that is removed from the text and audio modalities. Addressing these cross-modal vulnerabilities requires the development of adversarial multimodal re-identification models that can identify such risks and inspire solutions. Effective approaches will need to account for the interplay between modalities and ensure comprehensive anonymization across all channels of information.
@@
Current methods for generating multimodal synthetic data often lack integration with psychiatric knowledge, which limits their utility in mental health applications. Existing techniques typically generate synthetic data based solely on patient characteristics without incorporating cognitive models or therapeutic frameworks. Enhancing these methods with CBT-based models, akin to Patient-$\psi$ [28], could significantly improve the quality and relevance of the generated data. Additionally, most approaches rely on a sequential process, where synthetic text is generated first and then converted into audio and video using text-to-speech and video synthesis models. This pipeline can introduce inconsistencies and reduce authenticity. The development of multimodal synthetic data generators capable of producing therapy videos directly through advanced multimodal LLMs [118] represents a critical next step. Such systems could generate more cohesive and realistic synthetic data that better supports mental health research and applications.
@@
While privacy-utility evaluations have been explored for individual data modalities, there is a significant gap in understanding the trade-offs for multimodal data and models. The integration of cross-modal features may introduce unique vulnerabilities that require specialized evaluation frameworks. Additionally, comparative studies on the privacy-utility trade-offs between anonymized and synthetic data have yet to be conducted. Another challenge lies in addressing demographic-specific limitations. Some data anonymization methods perform poorly for certain demographic groups, such as those with distinct facial features [41]. Similarly, differential privacy-based training methods often exacerbate fairness issues, amplifying biases against underrepresented populations [117]. To tackle these challenges, privacy-utility evaluations must be conducted on diverse mental health datasets representing different demographics, cultures, and languages. This will ensure that privacy-preserving methods are inclusive, equitable, and effective across varied contexts.
@@ Conclusion
This paper highlights significant challenges of training and deploying AI models for real-world mental health diagnosis due to the sensitive nature of mental health data and the risks of private data leakage from trained models. To address these challenges, we examined key solutions, including data anonymization pipelines to remove PII, voice and face anonymization in therapy recordings, methods for generating synthetic data that replicate real-world scenarios without exposing real individuals, and differential privacy-based approaches for privacy-aware model training. Additionally, we detailed evaluation frameworks to assess the privacy and utility trade-offs of these methods, ensuring they maintain clinical relevance while safeguarding patient confidentiality. We proposed a comprehensive pipeline for developing privacy-aware mental health AI models, encompassing data collection, anonymization, synthetic data generation, privacy-utility evaluations, and privacy-aware training. This workflow aims to balance privacy protection with the utility required for effective mental health diagnosis and therapy assistance. Finally, we identified research prospects, such as advancing multimodal data anonymization techniques to address cross-modal vulnerabilities, improving synthetic data generation by integrating psychiatric knowledge and multimodal capabilities, and establishing robust evaluation frameworks for diverse demographics and cultures. These advancements will lay the groundwork for deploying privacy-preserving mental health AI systems in clinical settings, enabling better access to therapy while upholding the highest standards of data privacy and security.
@@
I.G., A.M. and T.C. contributed to conceptualizing the manuscript. A.M. led the effort of writing the initial draft of the manuscript. I.G., A.M. and T.C. finalized the manuscript. Tc. and I.G. supervised the project.
@@
This research work has been funded by the German Federal Ministry of Education and Research and the Hessian Ministry of Higher Education, Research, Science and the Arts within their joint support of the National Research Center for Applied Cybersecurity ATHENE. This work has also been funded by the LOEWE Distinguished Chair “Ubiquitous Knowledge Processing”, LOEWE initiative, Hesse, Germany (Grant Number: LOEWE/4a//519/05/00.002(0002)/81). T.C. acknowledges the support of Tower Research Capital Markets toward using machine learning for social good and Rajiv Khemani Young Faculty Chair Professorship in Artificial Intelligence.

### v2

@@ Introduction
Synthetic data must be rigorously tested for privacy vulnerabilities. Standard assessments should measure robustness against membership inference and attribute inference attacks [122, 51, 50, 49], as these attacks can reveal whether a specific individual’s data was part of the training set. The challenge is even greater in mental health datasets, where data is not only sparse but also highly diverse, making models more prone to overfitting on outliers. Thus, privacy gain [51] and outlier similarity [50] should be key evaluation criteria.
Beyond individual data points, synthetic datasets must also be scrutinized for direct content reproduction. Measuring reproduction rate [50] helps detect whether any real therapy sessions have been copied verbatim into the generated dataset. Additional privacy metrics like memorization coefficients [123], $\epsilon$-identifiability [124], and distance-based metrics [50, 49] should be employed to ensure a more comprehensive evaluation. Without these safeguards, synthetic data may introduce more risks than solutions.
@@
Synthetic data quality should be assessed through faithfulness (similarity to real-world data) and diversity (lexical, semantic, and topic variation) [105, 53]. This should be complemented by mental health professionals performing expert evaluations on aspects like naturalness, empathy, helpfulness, professionalism, comprehensiveness, authenticity, and safety [33, 54]. For multimodal synthetic data, metrics like coherence, contextual plausibility, overall naturalness of generated speech and gesture should also be measured [56, 57]. For large-scale evaluations, LLMs provide a scalable alternative to rate generated data on the same aspects as expert evaluations [33, 54]. For multimodal evaluations, multimodal LLMs [125] can be employed to assess the integration of speech, text, and video.
Ultimately, the best test of synthetic data utility lies in its real-world applicability. A crucial final evaluation step is training a multimodal model on the generated data and measuring its performance on a downstream mental health diagnosis task. Without strong predictive utility, privacy alone is insufficient.
@@
To ensure privacy in mental health AI, models must be rigorously tested for information leakage. Language models should be evaluated using canary insertion and membership inference attacks [39], while embeddings must be assessed for leakage of sensitive attributes (like location, age, gender, or identity) [40]. For models trained with DP-SGD, the privacy guarantee is quantified by the $\epsilon$-value [38] (which determines the distance within which errors are considered to be zero in Stochastic Gradient Descent). In FL, evaluations must account for gradient and weight leakage and FL specific attacks [116]. Without such targeted evaluations, privacy-preserving claims remain unverified and models may remain vulnerable to inference threats.
@@
Figure 2: Proposed pipeline for developing privacy-aware mental health AI. Our proposed pipeline for data collection and model training enables the development of privacy-aware mental health AI models, integrating safeguards to balance privacy with clinical utility.
@@
Developing privacy-preserving mental health AI systems begins with careful and ethically sound data collection. Therapy sessions must be recorded only with explicit, informed patient consent, detailing anonymization, secure storage, and exclusive research use. Approval from institutional ethics committees is essential. Audio should be captured using dual-channel recorders to separate therapist and patient voices, while video should focus on patients’ full facial and postural expressions for behavioral analysis. Transcription should be conducted locally using researchers or offline ASR tools to avoid third-party data exposure.
@@
Once collected, data must either be anonymized or replaced with synthetic data, depending on dataset scale and diversity. Small or demographically narrow datasets may require synthetic data augmentation to ensure sufficient coverage for training generalizable models. Larger, diverse datasets may benefit from direct anonymization. For anonymizing text transcripts, LLMs can be employed to detect and redact PII [28]. However, proprietary models cannot be used due to privacy issues. Thus, open source LLMs should be fine-tuned on augmented long-context conversational datasets for improved generalization [85, 86, 27]. Audio anonymization can be achieved by selectively replacing PII-bearing audio segments through either synthesizing matching audio segments through splicing in matching audio fragments [29] or emotion-aware voice synthesis [127] and voice conversion [62], preserving speech naturalness. Multi-speaker anonymization methods [52] should be used to maintain conversational dynamics while obscuring identity. For video, face anonymization tools such as FIVA [100] should be employed, though body language must also be obfuscated to prevent gender or identity leakage. Alternatively, synthetic datasets may be generated using multimodal LLMs [125] via role-playing as therapists and patients [53], zero-shot, or few-shot prompting [34], ensuring no direct mapping to real individuals.
@@
Both anonymized and synthetic data must be evaluated for privacy-utility trade-offs. Privacy evaluations should include re-identification tests using adversarial models [45], other related datasets [58] and measuring the effectiveness of techniques like multi-speaker anonymization through metrics such as EER [30, 31, 91, 92, 93] and FAR [52] in speaker verification. Voice anonymization should also be evaluated for unlinkability [93, 47] and robustness against attacks [30, 31, 91]. For video recordings, the privacy risks can be assessed using face verification systems and face reconstruction attacks[100] to determine the degree of obfuscation. Synthetic data must be rigorously tested against membership inference attacks and attribute inference attacks [122, 51, 50, 49]. Outlier leakage is another critical concern, especially in mental health datasets, where outliers are more prevalent due to the diversity and small size of participant groups. Metrics such as privacy gain [51], outlier similarity [50], and reproduction rate [50] are effective for evaluating these risks. These tests should be applied across languages, demographics, and cultures, and must incorporate cross-modal privacy assessments to detect information leakage through combination of modalities. The empirical privacy measures should also be accompanied by theoretical guarantees similar to $\epsilon$-value in DP. Utility evaluation should assess the usefulness of the data for mental health diagnosis tasks, focusing on information preservation [121, 45], intonation preservation [30], conversational diversity [30, 91], naturalness [30, 52], and emotional feature retention [31, 99]. LLMs can be leveraged to automatically evaluate the utility of synthetically generated data on dimensions such as naturalness, empathy, helpfulness, professionalism, comprehensiveness, authenticity, and safety [33, 54] or psychological measures like the working alliance inventory [53].
@@
During training, privacy-preserving methods like DP are critical. Mental health models either consist of models fine-tuned with mental health data or trained fusion layers to combine features extracted from various modalities using pre-trained models. To ensure privacy, these models should be fine-tuned with DP methods [42, 41], and fusion layers should be trained using DP-SGD [38]. In multi-institutional settings, federated learning combined with local differential privacy can offer stronger protection [116, 117].
@@
While Miao et al. [52] provided a benchmark for MSA, they assume weak attack models where the attacker does not have knowledge about the used anonymization scheme or cannot untangle speakers in overlapping segments. In reality, attackers may possess knowledge of the anonymization strategies or have the ability to separate overlapping speech segments, posing severe privacy risks in therapy settings. Consequently, MSA methods must be evaluated under stronger threat models and MSA should be enhanced to ensure resilience against sophisticated attacks.
@@
Video anonymization techniques also remain limited. Even after face obfuscation, models often leak sensitive attributes such as gender and age [32]. These attributes can be used to identify individuals especially in small mental health datasets with limited demographic representation. While recording the body language of patients could help in mental health diagnosis, it can also reveal the gender of the patient if only face anonymization is performed [101]. Moreover, current methods are also prone to demographic unfairness [48]. Thus it is essential to develop fair and improved video anonymization techniques that can prevent leakage of private information like age and gender.
@@
While we discuss various data anonymization processes for text, audio and video modalities, most of them do not provide any theoretical guarantees like DP provides in privacy-aware model training. In text modality, word-level or sentence-level perturbations through DP provide theoretical guarantees[128]. However, they significantly reduce the utility of the text [128], necessitating better anonymization techniques with privacy guarantees for text. For audio, DP guarantees exist only in single-speaker anonymization settings [93], whereas therapy sessions involve multiple speakers. Extending DP-based anonymization to multi-speaker contexts remains an open challenge. In video, existing DP methods for images [103] and video object indistinguishability [102] have not been adapted for privacy-preserving facial de-identification in therapy footage, calling for task-specific DP-based video anonymization techniques with validated utility and fairness. In synthetic data generation DP-based methods for theoretical guarantees have been explored for text [106, 107], tabular data [129, 130], multimodal tabular and 3D image data [131] generation and with FL [132]. However, no DP-based methods have been developed for multimodal therapy session generation.
@@
Multimodal data introduces unique privacy risks. Current anonymization methods treat each modality in isolation, but information can leak across modalities. For example, lip movements in video can reveal PII removed from text or audio. Addressing such vulnerabilities will require the development of multimodal adversarial re-identification models capable of identifying cross-modal risks and guiding the design of holistic anonymization techniques. This will require fundamentally new algorithms that jointly anonymize data across all modalities while preserving clinically relevant features.
@@
Current multimodal synthetic data generation methods often lack psychiatric grounding, reducing their utility in mental health applications. Most rely solely on patient characteristics without integrating therapeutic frameworks like Cognitive Behavioral Therapy (CBT). Incorporating such models, akin to Patient-$\psi$[35], could enhance clinical relevance of the generated data. Additionally, existing approaches follow sequential pipelines: text generation followed by audio and video synthesis leading to inconsistencies and diminished authenticity. Leveraging advanced multimodal LLMs [125] to generate cohesive, end-to-end therapy sessions provides a promising direction for improvement. However, current models struggle with coherence and memory in long multi-turn conversations [108], necessitating the development of long-context multimodal LLMs tailored to therapy data.
@@
Despite progress in modality-specific privacy-utility evaluations, assessments for multimodal data remain underdeveloped. Cross-modal interactions may introduce novel vulnerabilities that require specialized evaluation frameworks. Additionally, comparative studies on the privacy-utility trade-offs between anonymized and synthetic data have yet to be conducted. Furthermore, privacy-preserving techniques often underperform across demographic groups, particularly those with distinctive facial traits [48], and differential privacy methods may exacerbate these fairness issues, amplifying biases against underrepresented populations [126]. Addressing these gaps requires evaluation must be conducted on datasets that reflect diverse demographics, languages, and cultures. This will ensure that privacy-preserving methods are inclusive, equitable, and effective across varied contexts.
@@ Conclusion
This paper outlines key challenges in building AI systems for mental health diagnosis, emphasizing the privacy risks inherent in handling sensitive data. We critically reviewed core strategies to mitigate these risks, including text, voice, and face anonymization; synthetic data generation that replicates real-world interactions without revealing personal information; and privacy-preserving model training using differential privacy. Additionally, we detailed evaluation frameworks to assess the privacy and utility trade-offs of these methods to ensure clinical relevance is maintained while safeguarding confidentiality. Building on this, we proposed a comprehensive pipeline for developing privacy-aware mental health AI models, encompassing data collection, anonymization, synthetic data generation, privacy-utility evaluations, and privacy-aware training. This workflow aims to balance privacy protection with the utility required for effective mental health diagnosis and therapy assistance. Looking ahead, we highlight critical research directions: developing anonymization techniques that address cross-modal privacy risks, generating clinically grounded multimodal synthetic data, and establishing inclusive evaluation protocols across demographics and cultures. These advances are essential for deploying trustworthy, privacy-preserving AI systems that can expand access to mental health care while upholding the highest standards of data privacy and security.
@@
I.G., A.M. and T.C. contributed to conceptualizing the manuscript. A.M. led the effort of writing the initial draft of the manuscript. I.G., A.M. and T.C. finalized the manuscript. T.C. and I.G. supervised the project.
@@
This research work has been funded by the German Federal Ministry of Education and Research and the Hessian Ministry of Higher Education, Research, Science and the Arts within their joint support of the National Research Center for Applied Cybersecurity ATHENE. This work has also been funded by the LOEWE Distinguished Chair “Ubiquitous Knowledge Processing”, LOEWE initiative, Hesse, Germany (Grant Number: LOEWE/4a//519/05/00.002(0002)/81). This research was further supported by the DYNAMIC center, which is funded by the LOEWE program of the Hessian Ministry of Science and Arts (Grant Number: LOEWE/1/16/519/03/09.001(0009)/98). T.C. acknowledges travel support from the European Union’s Horizon 2020 research and innovation programme under Grant Agreement No 951847 and Rajiv Khemani Young Faculty Chair Professorship in Artificial Intelligence. T.C also acknowledges the support of the Alexander von Humboldt Foundation through a Humboldt Research Fellowship for Experienced Researchers.